# Comparison of Autoencoder Encodings for ECG Representation in Downstream Prediction Tasks

### v1

@@ Abstract
The electrocardiogram (ECG) is an inexpensive and widely available tool for cardiovascular assessment. Despite its standardized format and small file size, the high complexity and inter-individual variability of ECG signals (typically a 60,000-size vector) make it challenging to use in deep learning models, especially when only small datasets are available. This study addresses these challenges by exploring feature generation methods from representative beat ECGs, focusing on Principal Component Analysis (PCA) and Autoencoders to reduce data complexity. We introduce three novel Variational Autoencoder (VAE) variants—Stochastic Autoencoder (SAE), Annealed -VAE (A-VAE), and cyclical -VAE (C-VAE)—and compare their effectiveness in maintaining signal fidelity and enhancing downstream prediction tasks. The A-VAE achieved superior signal reconstruction, reducing the mean absolute error (MAE) to 15.73.2V, which is at the level of signal noise. Moreover, the SAE encodings, when combined with ECG summary features, improved the prediction of reduced Left Ventricular Ejection Fraction (LVEF), achieving an area under the receiver operating characteristic curve (AUROC) of 0.901. This performance nearly matches the 0.910 AUROC of state-of-the-art CNN models but requires significantly less data and computational resources. Our findings demonstrate that these VAE encodings are not only effective in simplifying ECG data but also provide a practical solution for applying deep learning in contexts with limited-scale labeled training data.
@@ III-E Downstream Clinically-relevant Prediction
To make comparisons, we used reduced LVEF as the clinically-relevant downstream prediction. LVEF is the percent of blood that is pumped by the left ventricle with each contraction. We had 303,265 ECGs on 105,370 patients paired with a unique echocardiogram within 45 days showing reduced LVEF (35%) in 14.1% [18][19].

### v2

@@ Abstract
The electrocardiogram (ECG) is an inexpensive and widely available tool for cardiovascular assessment. Despite its standardized format and small file size, the high complexity and inter-individual variability of ECG signals (typically a 120,000-size vector at 1000hz) make it challenging to use in deep learning models, especially when only small datasets are available. This study addresses these challenges by exploring feature generation methods from representative beat ECGs, focusing on Principal Component Analysis (PCA) and Autoencoders to reduce data complexity. We introduce three novel Variational Autoencoder (VAE) variants—Stochastic Autoencoder (SAE), Annealed -VAE (A-VAE), and cyclical -VAE (C-VAE)—and compare their effectiveness in maintaining signal fidelity and enhancing downstream prediction tasks. The A-VAE achieved superior signal reconstruction, reducing the mean absolute error (MAE) to 15.73.2V, which is at the level of signal noise. Moreover, the SAE encodings, when combined with ECG summary features, improved the prediction of reduced Left Ventricular Ejection Fraction (LVEF), achieving an area under the receiver operating characteristic curve (AUROC) of 0.901. This performance nearly matches the 0.910 AUROC of state-of-the-art CNN models but requires significantly less data and computational resources. Our findings demonstrate that these VAE encodings are not only effective in simplifying ECG data but also provide a practical solution for applying deep learning in contexts with limited-scale labeled training data.
@@ II-A Data
IRB Approval: This project was performed under the approval as an expedited review by the Institutional Review Board at The University of Kansas Medical Center (STUDY00160252). The ECGs for this project encompassed all clinically acquired ECGs at The University of Kansas Health System (2008-2022). ECGs were linked to variables obtained on clinical echocardiographic evaluations, when available, via Healthcare Enterprise Repository for Ontological Narration (HERON) linked by medical record numbers[7][8].
@@ III-E Downstream Clinically-relevant Prediction
To make comparisons, we used reduced LVEF as the clinically-relevant downstream prediction. LVEF is the percent of blood that is pumped by the left ventricle with each contraction. We had 303,265 ECGs on 105,370 patients paired with a unique echocardiogram within 45 days showing reduced LVEF (35%) in 14.1%.
@@ III-F VAE Performance
III-G Discussion
Our study presents significant advancements in applying autoencoder techniques, particularly our novel VAE variants, to ECG data analysis in clinical diagnostics. By effectively reducing high-dimensional ECG signals to a compact set of latent variables without requiring extensive datasets, we address a critical challenge in utilizing ECG data in machine learning models. Our novel VAE variants, particularly the A-VAE and SAE, demonstrate superior performance in preserving essential morphological features of the ECG by capturing non-linear relationships within the data, which is crucial for accurate clinical interpretations. Compared to traditional methods like PCA, our models better retain clinically relevant information, as evidenced by the improved prediction of complex ECG features such as bundle branch blocks and reduced LVEF. This enhanced capability leads to improved diagnostic accuracy, which is particularly impactful for conditions like heart failure where early detection can significantly alter patient outcomes.
By accommodating the inherent variability of ECG signals among individuals—through the SAE’s focus on reconstruction quality over latent space regularization—our approach allows for more personalized and accurate assessments, aligning with the move towards personalized medicine. Additionally, enabling robust predictive models in environments with limited datasets contributes to more equitable healthcare delivery by providing advanced diagnostic capabilities across under-represented and diverse populations. This extends the utility of VAEs in ECG analysis beyond interpretability to practical predictive performance, offering a pathway to integrate these models into existing clinical workflows even with smaller datasets.